A gas has a temperature of 14 °C, and a volume of 4.5 L. If the temperature is raised to 29 °C and the pressure is not changed, what is the new volume of the gas? Round to the nearest tenth. Don't forget the units.
1
Expert's answer
2012-12-24T11:05:17-0500
we have isobaric process here. Hence
V1/T1 = V2/T2

where T1=14 C = (273+14) K = 287 K;
T2=29 C = (273+29) K = 302 K
V1=4.5L

V2= V1 * T2/T1 = 4.5L * 302/287 = 4.7 L